Solve the momentum for each of the following moving objects: A. a 0.50kg basketball that is thrown upward with a velocity of 300 m/s. B. A 2000 kg light railway cart moving in the direction of south with a velocity of 10m/s
1
Expert's answer
2021-10-25T15:38:22-0400

A)

"p=mv=0.50\ kg\times300\ \dfrac{m}{s}=150\ \dfrac{kg\times m}{s}."


B)

"p=mv=2000\ kg\times10\ \dfrac{m}{s}=20000\ \dfrac{kg\times m}{s}."